In an auditorium there are 4a rows, and each row can accommodate 5yz benches. If x students can be accommodated in each bench, then how many students can sit in the auditorium?
Please help me with this problem I have tried and don't understand it. Please Help Me

Found 2 solutions by edjones, stanbon:
Answer by edjones(7311)

Turn this into something you can understand
If there are 4 rows of 5 benches with 7 students to a bench how many students can sit?
4*5*7=140
We multiplied all three.
So, 4a*5yz*x=20axyz Ans.
.
Ed

Answer by stanbon(48535)

In an auditorium there are 4a rows, and each row can accommodate 5yz benches. If x students can be accommodated in each bench, then how many students can sit in the auditorium?
---------------------------
Total # of benches: 4a*5yz = 20ayz
--------------------
Total # of places for students: x*20ayz = 20axyz
=================
Cheers,
Stan H.